### v1

# Scalable and Cost-Efficient de Novo Template-Based Molecular Generation

Abstract
Template-based molecular generation offers a promising avenue for drug design by ensuring generated compounds are synthetically accessible through predefined reaction templates and building blocks. In this work, we tackle three core challenges in template-based GFlowNets: (1) minimizing synthesis cost, (2) scaling to large building block libraries, and (3) effectively utilizing small fragment sets. We propose Recursive Cost Guidance, a backward policy framework that employs auxiliary machine learning models to approximate synthesis cost and viability. This guidance steers generation toward low-cost synthesis pathways, significantly enhancing cost-efficiency, molecular diversity, and quality, especially when paired with an Exploitation Penalty that balances the trade-off between exploration and exploitation. To enhance performance in smaller building block libraries, we develop a Dynamic Library mechanism that reuses intermediate high-reward states to construct full synthesis trees. Our approach establishes state-of-the-art results in template-based molecular generation.
1 Introduction
Generative models hold significant promise for accelerating drug discovery by enabling direct sampling of molecules with desired properties, vastly expanding the accessible chemical space. However, despite increasing interest from the machine learning community, their adoption in practical screening pipelines remains limited, primarily due to the poor synthesizability of the generated compounds. Recent research has begun to address this synthesizability bottleneck.
One prominent direction involves directly optimizing synthesizability, either by including synthesizability scores as part of the reward function [1], or by using retrosynthesis models to guide goal-conditioned generation [2]. However, such approaches remain fundamentally unreliable: heuristic metrics such as SAScore [3] often correlate poorly with true synthetic feasibility, while learned models such as RetroGNN [4] struggle with limited training data and poor generalization.
An increasingly popular and more principled alternative is template-based synthesis, where molecules are assembled from predefined building blocks through well-characterized reactions [5, 6]. This paradigm mirrors laboratory synthesis more closely and, with appropriate reaction templates, can provide an approximate guarantee of synthesizability. Recent work has embedded this formulation within Generative Flow Networks (GFlowNets) [7], which are well suited to molecular discovery due to their capacity for mode-seeking sampling and efficient exploration of combinatorial spaces [8, 9, 10].
Despite this progress, key limitations remain, particularly in terms of scalability and cost awareness. An ideal synthesis-aware generative model would explore the full space of synthesizable molecules while favoring those that are simpler and cheaper to make. Existing methods fall short: RGFN [8] employs a small, curated fragment library optimized for wet lab feasibility, but this restricts scalability. In contrast, RxnFlow [10] improves scalability through larger libraries and broader exploration, but ignores the cost of the synthesis, reducing its practical utility in downstream screening campaigns.
Figure 1: a) Recursive Cost Guidance employs machine learning models $\hat{c}_{B}^{S}$ and $\hat{c}_{B}^{D}$ to estimate the intractable synthesis cost and decomposability of precursor molecules, guiding the backward policy $P_{B}$ toward cheaper and viable intermediates. b) Forward policy is augmented with Exploitation Penalty to counter overexploitation induced by Synthesis Cost Guidance component of $P_{B}$. c) Dynamic Library gathers intermediate molecules with the highest expected reward and adds them to the building block library $M$, enabling full-tree synthesis.
In this paper, we propose SCENT (Scalable and Cost-Efficient de Novo Template-Based Molecular Generation) which addresses the aforementioned limitations of reaction-based GFlowNets. Our contributions can be summarized as follows (graphical summary in Figure 1):
  • We introduce Recursive Cost Guidance for the backward policy, which utilizes a machine learning model to approximate the recursive cost of backward transitions. From this, we derive two targeted strategies: Synthesis Cost Guidance, which reduces synthesis cost and improves diversity in large building block settings; and Decomposability Guidance, which enforces the validity of backward transitions by discouraging transitions through indecomposable molecules.
  • We analyze the exploitative behavior induced by Synthesis Cost Guidance and propose Exploitation Penalty, a simple yet effective regularizer that promotes exploration by penalizing repeated state-action pairs. This improves performance across all benchmarks.
  • We develop Dynamic Library mechanism that augments the initial set of building blocks with high-reward intermediates discovered during optimization. This mechanism enables full-tree synthesis and increases mode coverage in both low- and high-resource scenarios.
  • We benchmark recent reaction-based GFlowNets across three library sizes and three design tasks, demonstrating that SCENT significantly outperforms prior methods in terms of synthesis cost, molecular diversity, and the number of high-reward compounds discovered.
Our code is publicly available at https://github.com/koziarskilab/SCENT.
2 Related Works
Molecular Generation.
A broad range of machine learning approaches has been proposed for small molecule generation, varying in both molecular representations—such as graphs, SMILES, or 3D coordinates—and generative paradigms, including diffusion models [11, 12, 13, 14], flow matching [15, 16, 17], reinforcement learning [18, 19, 20], and large language models (LLMs) [21, 22, 23, 24]. Among these, Generative Flow Networks (GFlowNets) [7, 25, 26, 27, 28] have emerged as a compelling alternative to reinforcement learning for iterative molecular generation. Their ability to perform diverse, mode-covering sampling makes them particularly well suited for drug discovery.
Synthesis-Aware Molecular Generation.
Synthesis-aware molecular generation has gained momentum as a strategy to ensure that generated molecules are not only property-optimized but also synthetically accessible. These methods construct molecules from predefined building blocks using high-yield, well-characterized reactions. Recent GFlowNet-based approaches [9, 8, 10], along with reinforcement learning [5, 6], genetic algorithms [29], and autoregressive models [30] incorporate reaction-aware priors to produce more realistic candidates.
Backward Policy Design in GFlowNets.
In Generative Flow Networks (GFlowNets), the most common training objective is to match the forward policy $P_{F}$ and the backward policy $P_{B}$ to satisfy the trajectory balance condition [31]. While $P_{B}$ is typically learned jointly with $P_{F}$, it can also be optimized separately or fixed to influence the behavior of $P_{F}$. For example, Jang et al. [32] train $P_{B}$ via maximum likelihood over observed trajectories to induce exploitative behavior in $P_{F}$, whereas Mohammadpour et al. [33] use a fixed maximum-entropy $P_{B}$ to encourage exploration.
Guided Trajectory Balance (GTB) [34] is the most relevant prior work on backward policy guidance. GTB introduces a preference distribution over trajectories to a given terminal state and trains $P_{B}$ to match it. In contrast, we define trajectory preference through a recursive cost function, yielding a more interpretable and flexible formulation. We approximate this cost using an auxiliary model, which gives rise to our proposed Recursive Cost Guidance framework, which steers generation toward low-cost synthesis pathways while improving consistency with the underlying MDP.
SynFlowNet [9] also modifies $P_{B}$, using reinforcement learning to discourage transitions through nondecomposable intermediates. While it addresses similar challenges, our method provides a more general and principled solution by unifying synthesis cost reduction and MDP-consistent trajectory shaping through backward policy design.
Library Learning.
Library learning is a long-standing strategy for managing large search spaces in domains such as program induction and reinforcement learning. DreamCoder [35, 36, 37], for example, extracts reusable subroutines from sampled programs to build a library of primitives. Similarly, the Options framework [38] and macro-actions [39] learn temporally extended actions to facilitate exploration and improve credit assignment. More recently, ActionPiece [40] applied byte pair encoding (BPE) to compress frequent action subsequences in GFlowNet trajectories, augmenting the action space with learned macro-actions.
Our Dynamic Library builds on this foundation but shifts from action abstractions to state abstractions. Rather than compressing action sequences, we identify and cache high-reward intermediate states, which act as reusable substructures for constructing nonlinear synthesis trees. While conceptually similar to DreamCoder’s structural reuse, our approach operates in the space of synthesis pathways, a novel application of library learning principles to generative models for molecular design.
3 Method
In this section, we present SCENT and its core components: (1) Recursive Cost Guidance, a general backward policy framework that estimates recursive transition costs using a learned model; (1.1) Synthesis Cost Guidance, which biases generation toward low-cost synthesis pathways; (1.2) Decomposability Guidance, which discourages selection of indecomposable molecules; (2) Exploitation Penalty, which counterbalances the exploitative bias introduced by Synthesis Cost Guidance (see Section 4.4) and improves overall performance; and (3) Dynamic Building Block Library, which expands the building block set with high-reward intermediates to enable full-tree synthesis and trajectory compression. The complete pseudocode of SCENT’s training procedure is provided in Section A.1.
3.1 Preliminaries
Like other reaction-based GFlowNets [8, 9, 10], SCENT operates on a predefined set of reaction templates $R$ and building blocks $M$. Its forward policy $P_{F}$, adapted from RGFN (Section A.2), constructs molecules via sequential application of reactions. The backward policy $P_{B}$, in turn, decomposes a molecule by reversing these reactions. Specifically, given an intermediate molecule $m$, $P_{B}$ samples a tuple $(m^{\prime},M^{\prime},r)$ where $m^{\prime}$ is a precursor molecule, $M^{\prime}$ is a tuple of building blocks from $M$, and $r\in R$ is a reaction such that applying $r$ to $m^{\prime}$ and $M^{\prime}$ yields $m$. For clarity, we denote molecules $m$ and $m^{\prime}$ as states $s$ and $s^{\prime}$, and the pair $(M^{\prime},r)$ as an action $a^{\prime}$, so that the backward policy is defined as $P_{B}((s^{\prime},a^{\prime})\mid s)$.
The backward policy in SCENT is constructed using the Recursive Cost Guidance framework introduced in Section 3.2. It is composed of two sub-policies: the Synthesis Cost Guidance policy $P^{S}_{B}$ (Section 3.2.1) and the Decomposability Guidance policy $P^{D}_{B}$ (Section 3.2.2).
3.2 Recursive Cost Guidance
Recursive Cost Guidance is a new and general framework for backward policy guidance that employs a machine learning model to approximate the recursive costs of backward transitions. Let $c_{F}(\tau_{0:n})$ be a function that assigns a cost to a forward trajectory $\tau_{0:n}=(s_{0},a_{0},\dots,s_{n})$ starting from source state $s_{0}$:
$c_{F}(\tau_{0:n})=C(c_{F}(\tau_{0:n-1}),s_{n-1},a_{n-1},s_{n}),$ (1)
where $C$ defines the cost of transition from $s_{n-1}$ to $s_{n}$ via action $a_{n-1}$, possibly depending on the cumulative cost $c_{F}(\tau_{0:n-1})$. A natural instantiation of $c_{F}$ for template-based GFlowNets is the cost of the synthesis pathway, described in Section 3.2.1. To guide the backward policy $P_{B}$ toward trajectories with lower forward cost $c_{F}$, we define a backward cost function $c_{B}$ that represents the cost of the cheapest forward trajectory reaching a given state $s$:
$c_{B}(s)=\min_{\tau\in\mathcal{T}_{\rightarrow s}}c_{F}(\tau),$ (2)
where $\mathcal{T}_{\rightarrow s}$ is the set of all trajectories leading to state $s$. This can be reformulated into:
$c_{B}(s)=\min_{(s^{\prime},a^{\prime})\in SA^{\prime}}C(c_{B}(s^{\prime}),s^{\prime},a^{\prime},s),$ (3)
where $SA^{\prime}$ is a set of all parent state-action tuples. Since $c_{B}$ is defined recursively, exact computation is generally intractable. To address this, we estimate $c_{B}$ using a machine learning model $\hat{c}_{B}$, and define recursively guided backward policy $P_{B}$:
$P_{B}((s_{i}^{\prime},a_{i}^{\prime})|s)=\sigma_{i}^{|SA^{\prime}|}(\mathbf{l}),\;l_{i}=-\alpha C(\hat{c}_{B}(s^{\prime}_{i}),s^{\prime}_{i},a^{\prime}_{i},s),$ (4)
where $\alpha$ controls the greediness. This formulation avoids expensive recursive evaluations of $c_{B}$ by leveraging an inexpensive surrogate $\hat{c}_{B}$ (see Figure 3). To train $\hat{c}_{B}$, we maintain a fixed-size dataset $D$ containing current estimates of $c_{B}(s)$ denoted $\tilde{c}_{B}(s)$. After each iteration of forward policy training, we update these estimates using the encountered trajectories $\tau_{0:n}$ by applying: $\tilde{c}_{B}^{\prime}(s_{i})=\min(\tilde{c}_{B}^{\prime}(s_{i}),c_{F}(\tau_{0:i}))$. The model $\hat{c}_{B}$ is then fitted to $D$ using a dedicated cost loss $\mathcal{L}_{cost}$. In this work, we design two specific cost models — one for Synthesis Cost Guidance and another for Decomposability Guidance. The details of the models can be found in Section A.3.
Figure 2: Recursive Cost Guidance framework uses a cheap model $\hat{c}_{B}$ to approximate the intractable recursive cost $c_{B}$ of backward transitions.
Figure 3: Dynamic building block library is augmented using high-reward molecules (depicted in pink) that occur during sampling.
3.2.1 Synthesis Cost Guidance
Even though template-based approaches, if used with a curated set of reaction templates and fragments, can guarantee high synthesizability, the cost of synthesis of different molecules can vary widely depending on the cost of reactants and the yields of reactions. RGFN [8] approaches this issue by defining a highly curated set of low-cost fragments and high-yield reactions. However, this comes at the price of reducing molecular diversity and reachable chemical space. Instead, we propose defining a more extensive fragment library, but guiding the generation process toward cheaper molecules.
We estimate synthesis cost using stock prices of building blocks and reaction yields:
$C^{S}(x,s^{\prime},a^{\prime},s)=(x+c(a^{\prime}))y(s^{\prime},a^{\prime},s)^{-1},$ (5)
where $x$ represents either the accumulated cost of a synthesis pathway (in Equation 6) or the minimal cost to synthesize a molecule (in Equation 7). The term $c(a^{\prime})$ denotes the total cost of all reactants involved in the action $a^{\prime}$ ($a^{\prime}$ includes the building blocks $M^{\prime}$ and the reaction template $r$). The reaction yield $y(s^{\prime},a^{\prime},s)$ is the expected success rate of the reaction leading from reactants $\{s^{\prime}\}\cup M^{\prime}$ to the product $s$ via the template $r$. Details on how yields and building block costs are estimated are provided in Appendix C. While other factors, such as reaction time or solvent cost, may influence synthesis cost, they are either highly lab-specific or typically negligible. Consequently, our formulation of $C^{S}$ offers a robust and broadly applicable first-order approximation.
We integrate $C^{S}$ into the forward and backward cost formulations (Equation 1 and 3) as follows:
$c^{S}_{F}(\tau_{0:n})=C^{S}(c^{S}_{F}(\tau_{0:n-1}),s_{n-1},a_{n-1},s_{n}),$ (6)
$c^{S}_{B}(s)=\min_{(s^{\prime},a^{\prime})\in SA^{\prime}}C^{S}(c^{S}_{B}(s^{\prime}),s^{\prime},a^{\prime},s),$ (7)
As in Section 3.2, we use a machine learning model $\hat{c}^{S}_{B}$ to approximate the intractable $c^{S}_{B}$:
$P^{S}_{B}((s_{i}^{\prime},a_{i}^{\prime})|s)=\sigma_{i}^{|SA^{\prime}|}(\mathbf{l}),\;l_{i}=-\alpha_{S}C^{S}(\hat{c}^{S}_{B}(s^{\prime}_{i}),s^{\prime}_{i},a^{\prime}_{i},s),$ (8)
Further architectural and training details for $\hat{c}^{S}_{B}$ are provided in Section A.3. Our experiments show that the use of $P^{S}_{B}$ significantly reduces the cost of the proposed synthesis routes while dramatically increasing the number of distinct modes discovered (Section 4).
3.2.2 Decomposability Guidance
A core challenge in template-based GFlowNets is ensuring that the parent state $s^{\prime}$, sampled by the backward policy $P_{B}((s^{\prime},a^{\prime})\mid s)$, is recursively decomposable into building blocks. RGFN addresses this by explicitly verifying decomposability, but this incurs significant computational cost [8]. In contrast, Cretu et al. [9] propose two implicit strategies: (1) a reinforcement learning approach that assigns positive rewards to trajectories reaching the source state $s_{0}$ and negative rewards to those terminating in non-decomposable intermediates; and (2) a pessimistic backward policy trained via maximum likelihood over forward-sampled trajectories to implicitly favor valid paths.
In this work, we take a different approach by leveraging our Recursive Cost Guidance framework to address decomposability. Specifically, we define a backward cost function that penalizes invalid trajectories that do not begin at a source state:
$c^{D}_{B}(s)=\min_{(s^{\prime},a^{\prime})\in SA^{\prime}}c^{D}_{B}(s^{\prime}),$ (9)
with the base case $c^{D}_{B}(s)=0$ if $s$ is a source state and $SA^{\prime}=\emptyset$, and $c^{D}_{B}(s)=1$ otherwise. We then define the backward policy $P^{D}_{B}$, leveraging $\hat{c}^{D}_{B}$ model that approximates the intractable cost $c^{D}_{B}$:
$P^{D}_{B}((s_{i}^{\prime},a_{i}^{\prime})\mid s)=\sigma_{i}^{|SA^{\prime}|}(s),\quad l_{i}=-\alpha_{D}\hat{c}^{D}_{B}(s^{\prime}_{i}),$ (10)
Further architectural and training details for $\hat{c}^{D}_{B}$ are provided in Section A.3. Decomposability Guidance is a critical component of SCENT, substantially influencing overall performance (see Section G.1). Unless otherwise noted, all experiments in this paper assume that SCENT uses Decomposability Guidance. When Synthesis Cost Guidance is also enabled, SCENT combines the two by taking the product of their corresponding backward policies:
$P_{B}((s^{\prime},a^{\prime})\mid s)\propto P^{C}_{B}((s^{\prime},a^{\prime})\mid s)\cdot P^{D}_{B}((s^{\prime},a^{\prime})\mid s).$ (11)
3.3 Exploitation Penalty
As shown in Section 4.4, incorporating Synthesis Cost Guidance into SCENT induces more exploitative behavior. To counteract this tendency, we introduce a simple yet effective strategy that penalizes repeated action choices in a given state. Specifically, we augment the forward policy $P_{F}$ with an exploration-aware component $P_{E}$:
$P^{\prime}_{F}(a\mid s)\propto P_{F}(a\mid s)\cdot P_{E}(a\mid s),\quad P_{E}(a\mid s)=\frac{(n(s,a)+\epsilon)^{-\gamma}}{\sum_{a^{\prime}}(n(s,a^{\prime})+\epsilon)^{-\gamma}}.$ (12)
Here, $n(s,a)$ denotes the number of times action $a$ has been selected at state $s$, $\epsilon>0$ is a smoothing term, and $\gamma>0$ controls the strength of the penalty.
Unlike pseudo-count methods that rely on approximate density models [41, 42, 43], our method uses exact counts, avoiding estimation artifacts and simplifying implementation. Although this count-based approach requires explicit storage of visited state-action pairs—which could be problematic in large state spaces—we show in Section 4.4 that applying this penalty only during the initial training phase (e.g., the first 1,000 iterations) is sufficient to yield consistent exploration benefits. This keeps the memory footprint bounded and independent of the total number of training iterations.
3.4 Dynamic Building Block Library
Current template-based GFlowNets generate only left-leaning synthesis trees by drawing reactants from a fixed set of building blocks (Figure 3). We extend this setup with a Dynamic Building Block Library that augments the block set with high-reward intermediates discovered during training. These intermediates act as reusable subtrees, enabling the construction of richer, tree-structured synthesis pathways. This not only expands the space of synthesizable molecules but also allows trajectory compression (rewriting paths into compact trees), enhancing credit assignment and training efficiency.
Every $T$ training iterations, intermediate molecules encountered during the training are ranked by their expected utility $U(s)$:
$U(s)=\frac{1}{\lvert\mathcal{T}_{s}\rvert}\sum_{\tau\in\mathcal{T}_{s}}R(s_{\tau}),$ (13)
where $\mathcal{T}_{s}$ is the set of trajectories containing molecule $s$, and $R(s_{\tau})$ is the reward assigned by GFlowNet to the final state of the trajectory $\tau$. The top-$L$ intermediates, compatible with known reaction patterns, are added to the library. This procedure is update up to $N_{\text{add}}$ times. This approach significantly improves the performance of SCENT in all tested configurations (see Section 4).
4 Experiments
In Section 4.2, we compare SCENT against other template-based GFlowNets and show the contribution of individual SCENT’s components to overall performance. Section 4.3 analyzes the cost reduction mechanism behind the Synthesis Cost Guidance, Section 4.4 highlights its inherent exploitative behavior, motivating the Exploitation Penalty, and Section 4.5 demonstrates the scaling properties of the Synthesis Cost Guidance. Finally, Section 4.6 shows that SCENT displays improved synthesizability compared to other approaches.
4.1 Setup
Previously reported results for template-based GFlowNets are not directly comparable due to differences in template and building block sets. To ensure fairness, we introduce a unified benchmark with standardized templates and building blocks. We evaluated models in three settings: (1) SMALL: curated fragments (see Appendix C) for rapid synthesis or high-throughput screening, (2) MEDIUM: SynFlowNet’s templates with 64k Enamine building blocks, and (3) LARGE: like MEDIUM, but with 128k building blocks. These settings span practical use cases: SMALL focused on rapid synthesis and MEDIUM/LARGE supporting broader exploration requiring external procurement.
[…]
4.2 Evaluation of Template-Based Generative Models
[…]
Table 1: Template-based GFlowNets on the sEH proxy across SMALL, MEDIUM, and LARGE settings. SCENT (C) significantly reduces molecule cost and boosts scaffold and mode discovery in MEDIUM and LARGE. Dynamic Library (C+D) improves performance in all settings, while adding Exploitation Penalty (+P) yields the best overall results, making SCENT the strongest method.
model online discovery inference sampling
modes $>$ 8.0 $\uparrow$ scaff. $>$ 8.0 $\uparrow$ reward $\uparrow$ cost $>$ 8.0 $\downarrow$
SMALL RGFN 538 $\pm$21 5862 $\pm$354 7.34 $\pm$0.06 37.7 $\pm$2.0
SynFlowNet 10.0 $\pm$0.8 27.0 $\pm$8.6 6.89 $\pm$0.29 -
RxnFlow 11.3 $\pm$5.0 29.7 $\pm$12.7 4.18 $\pm$0.72 -
SCENT (w/o C) 510 $\pm$26 5413 $\pm$334 7.38 $\pm$0.02 37.1 $\pm$1.0
SCENT (C) 478 $\pm$11 5150 $\pm$87 7.43 $\pm$0.02 23.7 $\pm$4.0
SCENT (C+D) 596 $\pm$19 7148 $\pm$497 7.56 $\pm$0.05 19.7 $\pm$2.0
SCENT (C+P) 595 $\pm$20 6839 $\pm$360 7.42 $\pm$0.05 20.7 $\pm$2.6
SCENT (C+D+P) 715 $\pm$15 8768 $\pm$363 7.66 $\pm$0.03 20.7 $\pm$3.0
MEDIUM RGFN 4755 $\pm$541 10638 $\pm$918 7.08 $\pm$0.08 1268 $\pm$71
SynFlowNet 288 $\pm$28 299 $\pm$32 6.38 $\pm$0.03 1952 $\pm$548
RxnFlow 26.3 $\pm$3.9 27.3 $\pm$2.5 6.3 $\pm$0.02 -
SCENT (w/o C) 9310 $\pm$863 11478 $\pm$823 7.31 $\pm$0.09 1463 $\pm$62
SCENT (C) 17705 $\pm$4224 52340 $\pm$4303 7.74 $\pm$0.04 1163 $\pm$147
SCENT (C+D) 17241 $\pm$679 52607 $\pm$1352 7.71 $\pm$0.03 1141 $\pm$65
SCENT (C+P) 37714 $\pm$3430 90068 $\pm$9010 7.81 $\pm$0.04 1230 $\pm$121
SCENT (C+D+P) 32761 $\pm$1483 86081 $\pm$3171 7.75 $\pm$0.08 1117 $\pm$33
LARGE RGFN 5242 $\pm$123 6215 $\pm$195 7.09 $\pm$0.11 1725 $\pm$153
SynFlowNet 278 $\pm$282 385 $\pm$401 6.39 $\pm$0.19 -
RxnFlow 24.5 $\pm$2.5 25.0 $\pm$2.0 6.14 $\pm$0.08 -
SCENT (w/o C) 7171 $\pm$291 8767 $\pm$429 7.13 $\pm$0.12 1678 $\pm$63
SCENT (C) 12375 $\pm$264 36930 $\pm$5455 7.52 $\pm$0.09 1267 $\pm$159
SCENT (C+D) 17138 $\pm$1341 46443 $\pm$2080 7.68 $\pm$0.11 1256 $\pm$107
SCENT (C+P) 26367 $\pm$3193 46202 $\pm$10242 7.76 $\pm$0.03 1291 $\pm$66
SCENT (C+D+P) 24434 $\pm$4223 47816 $\pm$5538 7.78 $\pm$0.1 1399 $\pm$163
Introducing Synthesis Cost Guidance (C) into SCENT consistently reduces the synthesis cost across all settings while also improving the average reward. Importantly, it significantly increases the number of high-reward scaffolds in the MEDIUM and LARGE settings across all tasks. In the sEH task at the 8.0 threshold, Cost Guidance also substantially increases the number of high-reward modes; however, a slight decrease is observed at the 7.2 threshold. This indicates a shift toward more exploitative behavior, with the model focusing on high-reward regions. We analyze this trend in detail in Section 4.4, where we also motivate the introduction of the Exploitation Penalty.
Adding the Exploitation Penalty (P) to SCENT with Synthesis Cost Guidance (C+P) leads to consistent performance improvements across all settings and tasks, without sacrificing cost reduction. Notably, P recovers and surpasses the number of discovered modes at the 7.2 threshold, effectively mitigating the over-exploitation introduced by Synthesis Cost Guidance.
Incorporating the Dynamic Library (D) into SCENT with Synthesis Cost Guidance (C+D) yields further gains across all settings. These improvements are attributed to better credit assignment in compressed trajectories. The combination of P and D proves especially effective in the SMALL setting, suggesting complementary roles in improving exploration under constrained conditions.
Ultimately, SCENT with C+D+P in SMALL and C+P in MEDIUM and LARGE achieves the best overall performance in terms of diversity, quality, and cost of synthesis. These results highlight Cost Guidance’s dual role in reducing the cost of high-reward molecules and enabling effective scaling to larger building block libraries, as explored in Section 4.5.
4.3 Synthesis Cost Reduction Mechanism
We analyze how Cost Guidance reduces the average cost of the reaction path generated by $P_{F}$. As shown in Figure 4-a, Synthesis Cost Guidance shortens trajectory lengths, thereby decreasing overall path costs. Furthermore, Figure 4-b and -d demonstrate that Cost Guidance selects cheaper fragments and significantly limits the use of more expensive ones. Finally, Figure 4-c illustrates that the Synthesis Cost Guidance relies on a subset of fragments throughout training, which correlates with the frequent use of cheaper molecules. The corresponding analysis for the SMALL setting is provided in Appendix E. We perform an ablation study on different cost reduction methods in Section G.2.2, showing that Synthesis Cost Guidance consistently identifies more modes and higher-reward scaffolds, at lower costs.
Figure 4: Synthesis Cost Guidance reduces trajectory length (a), fragments cost (b), and reliance on expensive fragments (d), while concentrating on a smaller fragment subset (c). Results are smoothed over the last 100 iterations.
4.4 Synthesis Cost Guidance Exploitativeness Analysis
Figure 5: Revisit frequency of high-reward scaffolds increases sharply with Synthesis Cost Guidance (C), indicating greater exploitative behavior. Introducing the Exploitation Penalty (P) effectively reduces this revisit ratio.
The results in Section 4.2, together with the cost reduction mechanism detailed in Section 4.3, suggest that Synthesis Cost Guidance enhances the exploitative behavior of $P_{F}$. To validate this, we plot how often $P_{F}$ revisits previously discovered high-reward scaffolds during training in Figure 5. SCENT with Synthesis Cost Guidance (C) revisits the scaffolds drastically more often than SCENT without Synthesis Cost Guidance, confirming its increased exploitativeness. While this behavior promotes the discovery of high-reward scaffolds, it can reduce the diversity of discovered modes in certain settings, as observed in Appendix D. To address this trade-off, we introduced the Exploitation Penalty in Section 3.3, which drastically increases with the number of discovered modes and successfully limits scaffolds revisits in Figure 5.
4.5 Synthesis Cost Guidance Helps in Scaling
We conducted additional experiments to assess how the size of the building block library $M$ affects the performance gap between SCENT, SCENT without Cost Guidance, and SCENT with the Exploitation Penalty. As shown in Figure 6, Synthesis Cost Guidance consistently increases the number of discovered scaffolds across all library sizes (2k–128k). For mode discovery, the gap widens up to $M=64\text{k}$, then narrows at 128k.
Although initially counterintuitive that Cost Guidance improves mode-seeking, Section 4.3 suggests that it enables more efficient navigation of fragment space, reducing the number of fragments used. Previous work [9, 8] shows that reducing library size via random cropping can help, but the effect is modest compared to our gains. Similarly, removing high-cost fragments improves diversity only marginally (Section G.2.2). These findings support our conclusion that Cost Guidance learns to prioritize fragments that are both cost-efficient and reward-relevant.
Figure 6: Number of discovered a) scaffolds and b) modes as a function of the size of building block library $M$. The numbers are reported for the $3000$th training iteration.
4.6 Synthesibility-Oriented Metrics
We compare SCENT with leading GFlowNet baselines for molecular generation: RGFN and the fragment-based GFlowNet (FGFN) from Bengio et al. [7]. We also evaluate FGFN (SA), trained to jointly optimize the sEH proxy and the Synthetic Accessibility (SA) score. As shown in Figure 7, template-based models substantially improve synthesizability. This is most evident in the AiZynthFinder success rate [48], where FGFN fails completely and FGFN (SA) performs poorly relative to RGFN and SCENT. Notably, although FGFN (SA) achieves the lowest average SA score, it still struggles on AiZynthFinder, underscoring the limitations of heuristic synthesizability metrics. SCENT achieves the highest average reward and outperforms RGFN in synthesizability, which we attribute to its preference for shorter reaction pathways (Section 4.3), resulting in reduced molecular weight that correlates with the synthesibility metrics [49].
Figure 7: Synthesibility metrics for GFlowNet models. SCENT and RGFN were trained on SMALL.
5 Conclusion
In this paper, we presented SCENT, a template-based approach for molecular generation within the GFlowNet framework. SCENT introduces three new mechanisms that collectively enhance exploration of chemical space: (1) Resursive Cost Guidance, a general strategy for guiding GFlowNet’s backward policy, which we utilized to improve computational efficiency by removing costly parent state computation, and to focus sampling on molecules with low synthesis cost, (2) Dynamic Library building blocks, which improve exploration efficiency—especially with smaller fragment libraries—and enable the generation of non-linear synthesis trees, and (3) an Exploration Penalty, which avoids repeated visits to previously explored candidates. Through empirical evaluation, we demonstrated that combining these strategies leads to state-of-the-art results in template-based molecular generation across diverse experimental settings, from small-scale fragment libraries suitable for rapid, automated synthesis to large-scale libraries that create vast search spaces.
Appendix A SCENT Details
A.1 Pseudocode
Algorithm 1 SCENT training
  Input: building block library $M$, reactant patterns $R$, forward policy $P_{F}$, flags: $C$, $P$, and $D$.
  Initialize: sampling policy $P^{\prime}_{F}$, backward policy $P_{B}$.
   if $C$ then: $P_{B}=P^{D}_{B}\times P^{S}_{B}$, else: $P_{B}=P^{D}_{B}$.
   if $P$ then: $P^{\prime}_{F}=P_{F}\times P_{E}$, else: $P^{\prime}_{F}=P_{F}$ with $\epsilon$-greedy.
  for $i=1$ to $N$ do
     Sample forward trajectories $T_{F}$ from $P^{\prime}_{F}$.
     Sample replay trajectories $T_{B}$ from prioritized replay buffer using $P_{B}$.
     Update $P_{F}$ to minimize $\mathcal{L}_{TB}$ over $T=T_{F}\cup T_{B}$.
     Update Decomposability Guidance model $\hat{c}^{D}_{B}$.
     if $C$ then: Update Synthesis Cost Guidance model $\hat{c}^{S}_{B}$, else: do nothing.
     if $P$ then: Update Exploitation Penalty counter $n(s,a)$, else: do nothing.
     if $D$ then: Update Dynamic Library utility $U$ and every $T$ steps update $M$, else: do nothing.
  end for
A.2 Forward Policy
The backbone of SCENT’s $P_{F}$ is an adapted graph transformer model $f$ from [50]. It takes as input a partially constructed molecule $m$ (represented as a graph) and an optional reaction template $r$ (represented as a one-hot vector), and outputs an embedding $f(m,r)\in\mathbb{R}^{D}$. The building blocks from $M$ are embedded using a linear combination of the ECFP4 fingerprint $e(m)$ and a learnable embedding $v(m)$ uniquely associated with a given building block: $h(m)=We(m)+v(m)$.
A molecule (and a corresponding synthesis pathway) is composed in a sequence of stages that are further described in this section. All stages linearly transform the state embeddings obtained with a backbone model $f$ using weight matrices $W_{i}$.
Select an initial building block. In SCENT, the trajectory starts with a fragment $m$ sampled from $M$ with probability:
$p(m_{i})=\sigma_{i}^{|M|}(\mathbf{l}),\;l_{i}=\phi(W_{1}f(\emptyset,\emptyset))^{T}h(m_{i}),$ (14)
where $\phi$ is GELU activation function [51] and $\sigma^{k}$ is a standard softmax over the logits vector $\mathbf{l}\in\mathbb{R}^{k}$ of length $k$:
$\sigma_{i}^{k}(\mathbf{l})=\frac{\mathrm{exp}(l_{i})}{\sum_{j=1}^{k}\mathrm{exp}(l_{j})}.$
Select the reaction template. The next step is to sample a reaction $r_{i}\in R$ that can be applied to the molecule $m$:
$p(r_{i}|m)=\sigma_{i}^{|R|+1}(\mathbf{l}),\;l_{i}=\phi(W_{2}f(m,\emptyset))^{T}v(r_{i}),$ (15)
where $v(r_{i})$ is a learnable embedding for the $i$-th reaction. Choosing the stop action with the index $|R|+1$ ends the trajectory.
Select a set of reactants We want to find a building block $m_{i}\in M$ that will react with $m$ and reactants $M^{\prime}$ by the reaction template $r$. The probability of selecting $m_{i}$ is:
$p(m_{i}|m,r,M^{\prime})=\sigma_{i}^{|M|}(\mathbf{l}),$
$l_{i}=\phi(W_{3}g(m,r,M^{\prime}))^{T}h(m_{i})$
$g(m,r,M^{\prime})=f(m,r)+\sum_{m^{\prime}_{i}\in M^{\prime}}(h(m_{i}^{\prime})+v(i)),$
where $v(i)$ is a learnable positional embedding for the $i$-th position in the reaction template. If the reaction $r$ does not require any additional reactant, this step is skipped.
Perform the reaction and select one of the resulting molecules. In this step, we apply the reaction $r$ to a set of chosen fragments (it can be empty). We then sample the outcome molecule $m$ from a set of possible outcomes $O$:
$p(m_{i})=\sigma_{i}^{|O|}(\mathbf{l}),\;l_{i}=\phi(W_{4}(f(m_{i},\emptyset)))^{T}w,$ (16)
where $w\in\mathbb{R}^{D}$ is a learnable parameter.
A.3 Cost Prediction Model Training
We use two cost prediction models in SCENT: synthesis cost prediction model $\hat{c}_{B}^{S}$ and decomposability prediction model $\hat{c}_{B}^{D}$. The size of mini-batches, the dataset size $N$, and the negative sampling ratio were manually chosen to minimize the cost prediction validation losses gathered during the training.
A.3.1 Synthesis Cost Model
The cost prediction model $\hat{c}_{B}^{S}$ is a simple multi-layer perception on top of ECFP4 fingerprint $e(s)\in\mathbb{R}^{2048}$:
$\hat{c}(s)=W_{1}\phi(W_{2}e(s)),\;W_{1}\in\mathbb{R}^{1\times 128},\;W_{2}\in\mathbb{R}^{128\times 2048},$ (17)
where $\phi$ is the GELU activation function [51].
We train the model $\hat{c}_{B}^{S}$ using the procedure described in Section 3.2. Given an updated dataset with synthesis cost estimates $D$, we perform five mini-batch updates of $\hat{c}_{B}^{S}$ using mini-batches of size $1024$, the learning rate of $0.01$ and the mean square error loss $\mathcal{L}_{cost}=\sum_{c_{s}\in D^{\prime}}\lvert\lvert c_{s}-\hat{c}_{B}^{S}\rvert\rvert_{2}^{2}$, where $D^{\prime}$ is a mini-batch of $D$. To stabilize training, we standardize all costs in $D$ using the mean and variance of costs from the building block library $M$.
A.3.2 Decomposability Model
The decomposability model $\hat{c}_{B}^{D}$ is very similar to the synthesis cost model, except that it encodes the current number of reactions used in the molecule $s$. The "cost" in the context of decomposability is 00 if the given molecule can be decomposed into building blocks from $M$ (within some number of steps) and $1$ otherwise. In practice, we train the model $\hat{c}_{B}^{D}$ to predict the decomposability label $\in\{0,1\}$ using binary cross entropy $\mathcal{L}_{cost}$. The labels are gathered similarly to the synthesis cost values, using the trajectories encountered in the GFlowNet forward sampling. We update $\hat{d}(s)$ with five minibatches of size $1024$ and a learning rate of $0.005$. In addition, we ensure that at least $20\%$ of the samples of molecules for training $\hat{d}(s)$ are not decomposable.
Appendix B Training Details
All the models sampled 320,000 forward trajectories during the training in total in SMALL and MEDIUM settings and 256,000 in LARGE. All forward policies were trained using the Trajectory Balance objective [31] and their parameters were optimized using Adam [52]. All methods were trained using their built-in action embedding mechanisms and on a maximum number of reactions equal to $4$. Hyperparameters were chosen semi-manually or using small grid searches to maximize the number of high-reward modes in the SMALL setting in the sEH proxy.
B.1 SCENT
[…]
All SCENT instances were trained using Decomposability Guidance, either solely or with Synthesis Cost Guidance using Equation 11.
Decomposability Guidance
To predict the decomposability of the molecule, we use the model described in Section A.3. We set the cost temperature (Equation 10) to $\alpha_{D}=5$.
Synthesis Cost Guidance
The model for Synthesis Cost Guidance is described in Section A.3. We set the cost temperature to $\alpha=5$.
Dynamic Building Block Library
We update the dynamic building block library every $T=1000$ iterations for a maximum of $N_{\text{add}}=10$ additions (that is, we do this until the end of training). The number of molecules added to the library every time it is updated is $L=400$.
Exploitation Penalty
We set $\epsilon=3$ in Equation 12 and schedule $\gamma$ in two ways: 1) so that it linearly decays to zero after $N$ iterations, and 2) it grows with the trajectory length (each step increases it by some constant factor $\Delta\gamma$). We set $N=\infty$ for SCENT that uses Dynamic Library and $N=1000$ for the rest of the runs. The growing delta $\Delta\gamma=0.2$, while the initial $s_{0}$ temperature $\gamma_{0}=1.0$. Note that for JNK3 in the MEDIUM setting, we used $4500$ iterations for exploitation penalty.
B.2 RGFN
[…]
B.3 SynFlowNet
[…]
We trained SynFlowNet using $32$ forward trajectories and $8000$ iterations on LARGE setting, and $64$ forward trajectories and $5000$ iterations for SMALL and MEDIUM.
B.4 RxnFlow
[…]
Appendix C Reaction Yields and Fragments Costs Estimates
SMALL setting.
The reaction set for the SMALL setting is the extended reaction set from Koziarski et al. [8]. We added Bishler-Napieralski and Pictet-Spengler reactions, along with various aryl functionalizations, benzoxazole, benzimidazole, and benzothiazole formation reactions, Hantzsch thiazole synthesis, alkylation of aromatic nitrogen, and Williamson ether synthesis.
Building block prices for SMALL were obtained from the vendors online. If multiple vendors offered the same building block, only the cost of the cheapest option was considered. For compounds available in varying amounts, both the price per gram and the alternative size (e.g., 5, 10, 25 grams) were considered, and the smallest price per mmol was used. Compounds that were exclusively accessible in other forms, such as salts, related functional groups, or containing a protected functional group were used with the SMILES fitted to the corresponding reactions. The cost of a product includes the total cost of the building blocks used, while additional expenses such as solvents, catalysts, and reagents are not accounted for.
Reaction yield estimates for amide coupling, nucleophilic additions to isocyanates, Suzuki reaction, Buchwald-Hartwig coupling, Sonogashira coupling, and azide-alkyne cycloaddition were obtained by analyzing in-house data. For all other reactions, SMARTS templates were used to search for reaction substructure in Reaxys and SciFinder, and the obtained literature data was analyzed to produce yield estimates.
MEDIUM and LARGE setting.
We used Enamine building blocks subsets of size 64,000 for MEDIUM and 128,000 for LARGE settings. The Enamine REAL database was first filtered to retain only molecules with currently available cost information. Fragment selection was then performed randomly for LARGE setting, without any specific criteria, to ensure broad coverage of molecular space. The building blocks for MEDIUM setting were subsampled from LARGE one.
Appendix D Extended Comparison to Template-Based GFlowNets
D.1 Results on sEH with Reward Threshold = 7.2
Despite tuning the SynFlowNet [9] parameters (Appendix B), we could not obtain satisfying results. Importantly, our experimental setting differs from the one used in their paper in the following way: 1) we use up to 4 reactions, they use 2 or 3, 2) they use a 10k subset of Enamine, our MEDIUM setting contains 64k molecules, 3) we set the reward thresholds higher (for sEH we set 8.0 instead of 7.2), and 4) we train the model for fewer trajectories. In Table 2, we report the results for sEH using a threshold from the SynFlowNet paper. The number of modes obtained seems consistent with those reported in Cretu et al. [9].
Table 2: Evaluation of different template-based approaches using sEH proxy in SMALL, MEDIUM and LARGE settings, but using lower reward threshold ($>7.2$) equivalent to the one used in SynFlowNet paper [9].
model online discovery inference sampling
modes $>$ 7.2 $\uparrow$ scaff. $>$ 7.2 $\uparrow$ reward $\uparrow$ cost $>$ 7.2 $\downarrow$
SMALL RGFN 11320 $\pm$168 76742 $\pm$693 7.34 $\pm$0.06 35.3 $\pm$1.6
SynFlowNet 559 $\pm$217 4795 $\pm$2477 6.89 $\pm$0.29 22.6 $\pm$7.3
RxnFlow 472 $\pm$259 734 $\pm$270 4.18 $\pm$0.72 52.3 $\pm$5.9
SCENT (w/o C) 13862 $\pm$422 114522 $\pm$1558 7.38 $\pm$0.02 38.8 $\pm$1.0
SCENT (C) 11978 $\pm$197 108262 $\pm$1355 7.43 $\pm$0.02 23.5 $\pm$2.8
SCENT (C+D) 11877 $\pm$48 110522 $\pm$1632 7.56 $\pm$0.05 18.9 $\pm$2.8
SCENT (C+P) 14399 $\pm$1149 137443 $\pm$1850 7.42 $\pm$0.05 19.5 $\pm$1.3
SCENT (C+D+P) 14631 $\pm$206 137426 $\pm$1055 7.66 $\pm$0.03 20.9 $\pm$1.7
MEDIUM RGFN 45937 $\pm$3108 77876 $\pm$3078 7.08 $\pm$0.08 1386 $\pm$64
SynFlowNet 32310 $\pm$682 35481 $\pm$643 6.38 $\pm$0.03 1666 $\pm$109
RxnFlow 2458 $\pm$46 2613 $\pm$45 6.3 $\pm$0.02 1645 $\pm$56
SCENT (w/o C) 126409 $\pm$4001 141284 $\pm$2833 7.31 $\pm$0.09 1594 $\pm$59
SCENT (C) 108921 $\pm$17019 212402 $\pm$1361 7.74 $\pm$0.04 1151 $\pm$140
SCENT (C+D) 104702 $\pm$3536 213750 $\pm$1818 7.71 $\pm$0.03 1150 $\pm$60
SCENT (C+P) 158578 $\pm$3617 277811 $\pm$4383 7.81 $\pm$0.04 1230 $\pm$132
SCENT (C+D+P) 144224 $\pm$2329 274380 $\pm$1388 7.75 $\pm$0.08 1096 $\pm$29
LARGE RGFN 54255 $\pm$1573 59210 $\pm$1146 7.09 $\pm$0.11 1766 $\pm$50
SynFlowNet 20186 $\pm$13305 36656 $\pm$22657 6.39 $\pm$0.19 1545 $\pm$70
RxnFlow 1615 $\pm$60 1708 $\pm$67 6.14 $\pm$0.08 1800 $\pm$22
SCENT (w/o C) 95457 $\pm$1292 104789 $\pm$1684 7.13 $\pm$0.12 1776 $\pm$50
SCENT (C) 83729 $\pm$1157 157331 $\pm$1783 7.52 $\pm$0.09 1230 $\pm$159
SCENT (C+D) 90133 $\pm$3566 164962 $\pm$2931 7.68 $\pm$0.11 1267 $\pm$116
SCENT (C+P) 140654 $\pm$6121 198503 $\pm$5047 7.76 $\pm$0.03 1250 $\pm$61
SCENT (C+D+P) 136677 $\pm$3734 199216 $\pm$5986 7.78 $\pm$0.1 1362 $\pm$146
[…]
model online discovery inference sampling
modes $>$ 0.8 $\uparrow$ scaff. $>$ 0.8 $\uparrow$ reward $\uparrow$ cost $>$ 0.8 $\downarrow$
SMALL RGFN 285 $\pm$14 1870 $\pm$106 0.7 $\pm$0.01 42.3 $\pm$2.7
SynFlowNet 0.0 $\pm$0.0 0.0 $\pm$0.0 0.57 $\pm$0.01 -
RxnFlow 2.0 $\pm$0.0 2.0 $\pm$0.0 0.51 $\pm$0.05 -
SCENT (w/o C) 290 $\pm$17 1911 $\pm$84 0.7 $\pm$0.01 44.1 $\pm$5.1
SCENT (C) 287 $\pm$9 1856 $\pm$147 0.7 $\pm$0.0 34.9 $\pm$4.5
SCENT (C+D) 306 $\pm$14 2138 $\pm$25 0.73 $\pm$0.01 30.3 $\pm$1.3
SCENT (C+P) 294 $\pm$18 2000 $\pm$109 0.71 $\pm$0.0 30.0 $\pm$0.4
SCENT (C+D+P) 336 $\pm$6 2172 $\pm$92 0.72 $\pm$0.01 32.1 $\pm$3.1
MEDIUM RGFN 4037 $\pm$1278 5225 $\pm$1122 0.64 $\pm$0.01 1251 $\pm$211
SynFlowNet 681 $\pm$184 1015 $\pm$393 0.58 $\pm$0.02 1565 $\pm$415
RxnFlow 64.0 $\pm$14.8 88.0 $\pm$11.2 0.58 $\pm$0.0 1602 $\pm$59
SCENT (w/o C) 6430 $\pm$2427 10010 $\pm$2537 0.66 $\pm$0.03 1177 $\pm$187
SCENT (C) 5945 $\pm$4009 50435 $\pm$13360 0.74 $\pm$0.03 917 $\pm$200
SCENT (C+D) 9430 $\pm$4284 64721 $\pm$15868 0.77 $\pm$0.01 776 $\pm$247
SCENT (C+P) 8868 $\pm$1883 92813 $\pm$34174 0.77 $\pm$0.05 815 $\pm$56
SCENT (C+D+P) 11857 $\pm$4171 93761 $\pm$39462 0.82 $\pm$0.04 730 $\pm$184
LARGE RGFN 1748 $\pm$660 2056 $\pm$638 0.62 $\pm$0.02 1370 $\pm$176
SynFlowNet 402 $\pm$237 970 $\pm$687 0.53 $\pm$0.01 -
RxnFlow 31.7 $\pm$5.6 37.7 $\pm$5.0 0.55 $\pm$0.0 -
SCENT (w/o C) 3366 $\pm$3503 4873 $\pm$4494 0.6 $\pm$0.07 1842 $\pm$606
SCENT (C) 4295 $\pm$1468 28867 $\pm$2418 0.7 $\pm$0.01 973 $\pm$119
SCENT (C+D) 8238 $\pm$2716 37957 $\pm$3277 0.74 $\pm$0.01 693 $\pm$107
SCENT (C+P) 7890 $\pm$1869 26553 $\pm$2121 0.73 $\pm$0.01 931 $\pm$197
SCENT (C+D+P) 5574 $\pm$708 34595 $\pm$3254 0.73 $\pm$0.03 990 $\pm$151
Table 4: Evaluation of different template-based approaches using JNK3 proxy in SMALL, MEDIUM and LARGE settings. In all metrics, except "reward", we consider molecules with reward $>0.8$. We observe that SCENT (C) that uses Cost Guidance can greatly reduce the cost of the sampled molecules in SMALL and MEDIUM settings, while drastically increasing the number of scaffolds.
model online discovery inference sampling
modes $>$ 0.8 $\uparrow$ scaff. $>$ 0.8 $\uparrow$ reward $\uparrow$ cost $>$ 0.8 $\downarrow$
SMALL RGFN 307 $\pm$11 2852 $\pm$42 0.8 $\pm$0.0 11.6 $\pm$0.3
SynFlowNet 0.0 $\pm$0.0 0.0 $\pm$0.0 0.42 $\pm$0.08 -
RxnFlow 16.3 $\pm$1.9 70.3 $\pm$4.1 0.69 $\pm$0.0 12.2 $\pm$0.2
SCENT (w/o C) 271 $\pm$9 2363 $\pm$100 0.78 $\pm$0.01 16.8 $\pm$2.8
SCENT (C) 256 $\pm$11 2248 $\pm$138 0.8 $\pm$0.01 9.27 $\pm$0.5
SCENT (C+D) 293 $\pm$1 2771 $\pm$38 0.79 $\pm$0.0 8.82 $\pm$0.36
SCENT (C+P) 247 $\pm$12 2395 $\pm$191 0.8 $\pm$0.02 12.0 $\pm$3.4
SCENT (C+D+P) 303 $\pm$6 2765 $\pm$150 0.8 $\pm$0.0 9.45 $\pm$0.55
MEDIUM RGFN 737 $\pm$62 5685 $\pm$969 0.6 $\pm$0.01 1699 $\pm$83
SynFlowNet 0.33 $\pm$0.47 0.33 $\pm$0.47 0.5 $\pm$0.0 -
RxnFlow 4.0 $\pm$1.0 4.0 $\pm$1.0 0.51 $\pm$0.01 -
SCENT (w/o C) 711 $\pm$75 11383 $\pm$5011 0.62 $\pm$0.03 1599 $\pm$248
SCENT (C) 525 $\pm$61 51264 $\pm$23502 0.77 $\pm$0.02 1197 $\pm$320
SCENT (C+D) 423 $\pm$108 60061 $\pm$31036 0.81 $\pm$0.01 1063 $\pm$352
SCENT (C+P) 637 $\pm$32 91872 $\pm$5942 0.76 $\pm$0.01 869 $\pm$38
SCENT (C+D+P) 645 $\pm$330 89508 $\pm$53412 0.79 $\pm$0.06 1909 $\pm$395
LARGE RGFN 203 $\pm$7 305 $\pm$14 0.63 $\pm$0.02 1652 $\pm$59
SynFlowNet 5.0 $\pm$7.07 83.3 $\pm$117.9 0.51 $\pm$0.1 -
RxnFlow 3.0 $\pm$1.63 3.67 $\pm$2.49 0.5 $\pm$0.01 -
SCENT (w/o C) 430 $\pm$295 1817 $\pm$1219 0.58 $\pm$0.01 2454 $\pm$637
SCENT (C) 116 $\pm$44 3165 $\pm$2437 0.67 $\pm$0.02 1213 $\pm$186
SCENT (C+D) 113 $\pm$87 10412 $\pm$5619 0.71 $\pm$0.01 1403 $\pm$405
SCENT (C+P) 442 $\pm$325 15192 $\pm$9772 0.71 $\pm$0.04 1383 $\pm$126
SCENT (C+D+P) 472 $\pm$238 11616 $\pm$5744 0.71 $\pm$0.01 1278 $\pm$43
[…]
D.3 Results on JNK3
In Table 4 RGFN outperforms all SCENT variants in the smallest setting, consistently finding more diverse modes and high-reward scaffolds. While SCENT lags behind in terms of mode discovery, it compensates in terms of cost, since SCENT (C) samples the cheapest molecules. In the MEDIUM setting, SCENT (w/o C) outperforms the baseline methods and the fully enabled SCENT variants in terms of mode discovery, but lags behind in the number of high-reward scaffolds. That can be attributed to the fact that SCENT (C) samples are less diverse. However, SCENT (C+P) samples the cheapest molecules. Finally, in the largest setting, the fully enabled SCENT (C+P) finds the most number of modes and high-reward scaffolds, whereas SCENT (C) takes the spotlight in terms of cost, generating the cheapest molecules.
Appendix E Extended Synthesis Cost Guidance Analysis
In this section, we extend the analysis from Section 4.3 with results in the SMALL setting. Figure 8 shows that similarly to the MEDIUM setting, Synthesis Cost Guidance reduces the average length of the trajectories and the average price of the fragments used also in the SMALL setting. Interestingly, it does not significantly limit the usage of fragments, which is likely due to their shortage and relatively uniform cost. We also see that the reaction yield does not play a role in reducing the cost.
Figure 8: Cost reduction analysis for sEH proxy and SMALL setting. Cost Guidance decreases the average length of the sampled trajectory (a), and the average cost of used fragments (b) which directly reduces the synthesis path cost. Importantly, the average reaction yield (d) does not increase when using Cost Guidance, suggesting its low influence in reducing the synthesis costs. The plots are smoothened by averaging results from the last 100 iterations.
Appendix F Further Comparison to Baselines
In this section, we extend our comparison of SCENT to established baseline models from the literature: REINVENT [20] and Graph GA [53]. Figure 9 shows a comparative analysis of the molecular properties generated by each method, filtered through AIZynthFinder [48] to retain only retrosynthetically accessible compounds. SCENT achieves a reward comparable to REINVENT, while outperforming it in terms of diversity and novelty. Graph GA generates highly diverse molecules, but these are often challenging to synthesize, as indicated by its low support in the final sub-plot. Importantly, both SCENT and REINVENT lie on the Pareto front of the reward–novelty trade-off: SCENT is biased toward higher novelty, whereas REINVENT tends toward higher reward. LABEL:tab:synth_metrics_1 and LABEL:tab:synth_metrics_2 further support this comparison, confirming that SCENT generates diverse, high-quality, and synthetically accessible molecules.
Figure 9: Comparison of molecules sampled by SCENT, REINVENT, and Graph GA, filtered using AIZynthFinder to retain only synthetically accessible candidates. Diversity is measured as the average Tanimoto distance between a molecule and the rest of the sampled set, while novelty is defined as the average Tanimoto distance to the closest molecule in the ChEMBL database. SCENT was trained in the SMALL setting. We observe that it generates diverse, high-reward, and synthesizable molecules, and lies on the Pareto front of the novelty–reward trade-off. Tick marks in the final sub-plots indicate the mean values of the sampled distributions.
Table 5: Comparison of generative models. Diversity, Novelty, and AIZynth scores were computed on the top 100 highest-scoring molecules. Diversity is defined as the average Tanimoto distance between each molecule and the rest of the sampled set. Novelty refers to the average Tanimoto distance to the closest molecule in the ChEMBL database. AIZynth represents the percentage of molecules successfully retrosynthesized by the AIZynthFinder software. All results are averaged over three independent runs of training and evaluation. Template-based models were trained under the SMALL setting.
Setting Reward $\uparrow$ Diversity $\uparrow$ Novelty $\uparrow$ AIZynth $\uparrow$
Template-based RGFN 7.34 $\pm$0.06 0.77 $\pm$0.02 0.58 $\pm$0.01 0.647 $\pm$0.037
SynFlowNet 6.89 $\pm$0.29 0.74 $\pm$0.02 0.53 $\pm$0.01 0.247 $\pm$0.090
RxnFlow 6.30 $\pm$0.03 0.80 $\pm$0.00 0.53 $\pm$0.01 0.653 $\pm$0.021
SCENT (C+D+P) 7.66 $\pm$0.03 0.73 $\pm$0.01 0.59 $\pm$0.01 0.773 $\pm$0.066
Unconstrained Graph GA 7.30 $\pm$0.31 0.56 $\pm$0.05 0.62 $\pm$0.08 0.113 $\pm$0.160
FGFN 6.03 $\pm$0.10 0.85 $\pm$0.00 0.66 $\pm$0.00 0.010 $\pm$0.000
FGFN + SA 6.16 $\pm$0.20 0.80 $\pm$0.01 0.54 $\pm$0.02 0.453 $\pm$0.091
REINVENT 7.97 $\pm$0.07 0.59 $\pm$0.10 0.50 $\pm$0.05 0.800 $\pm$0.113
Table 6: Comparison of generative models on Quantitative Estimation of Drug-likeness (QED), weight, Synthetic Complexity (SC), and Synthetic Accessibility (SA) metrics. Template-based models were trained on SMALL setting.
Setting QED $\uparrow$ weight $\downarrow$ SC $\downarrow$ SA $\downarrow$
Template-based RGFN 0.244 $\pm$0.003 542 $\pm$7 4.84 $\pm$0.01 3.30 $\pm$0.03
SynFlowNet 0.454 $\pm$0.113 451 $\pm$76 4.42 $\pm$0.28 3.06 $\pm$0.11
RxnFlow 0.259 $\pm$0.007 603 $\pm$9 3.34 $\pm$0.04 4.73 $\pm$0.06
SCENT (C+D+P) 0.286 $\pm$0.004 519 $\pm$5 4.82 $\pm$0.01 3.24 $\pm$0.02
Unconstrained Graph GA 0.464 $\pm$0.033 367 $\pm$12 3.72 $\pm$0.23 3.55 $\pm$0.66
FGFN 0.197 $\pm$0.003 668 $\pm$5 4.88 $\pm$0.01 4.77 $\pm$0.06
FGFN + SA 0.186 $\pm$0.004 621 $\pm$10 4.77 $\pm$0.04 3.42 $\pm$0.12
REINVENT 0.188 $\pm$0.088 655 $\pm$179 4.64 $\pm$0.07 2.76 $\pm$0.12
Appendix G Ablations
G.1 Decomposability Guidance Ablation
Decomposability Guidance, described in Section 3.2.2, is an integral part of our SCENT model. Leveraging a machine learning model, it predicts which molecules can be further decomposed in the backward transition (retrosynthesized) and discards the transitions that would otherwise lead to invalid states. Table 7 shows that this component is crucial for the performance of SCENT.
Table 7: Effect of Decomposability guidance on online discovery and inference in SCENT.
model online discovery inference sampling
modes $>$ 8.0 $\uparrow$ scaff. $>$ 8.0 $\uparrow$ reward $\uparrow$ cost $>$ 8.0 $\downarrow$
SMALL w/o Decomposability 309 $\pm$4 2794 $\pm$77 7.3 $\pm$0.03 52.2 $\pm$4.9
w/ Decomposability 510 $\pm$26 5413 $\pm$334 7.38 $\pm$0.02 37.1 $\pm$1.0
MEDIUM w/o Decomposability 6989 $\pm$303 8840 $\pm$652 7.15 $\pm$0.05 1549 $\pm$81
w/ Decomposability 9310 $\pm$863 11478 $\pm$823 7.31 $\pm$0.09 1463 $\pm$62
G.2 Synthesibility Cost Guidance Ablations
G.2.1 Cost Prediction Model Ablation
In this section, we demonstrate the importance of accurately estimating the true cost $c$ using the machine learning model $c^{\prime}$ for the optimal performance of SCENT. We compare our model with a low-quality cost estimator that outputs a constant value for every input molecule, where this constant is the mean of all the building blocks in the given scenario. It is important to note that we only modify the recursive component of the $C^{S}$ function from Equation 5, keeping the fragment costs unchanged. As shown in Table 8, the use of a constant value can reduce costs and improve SCENT’s performance in MEDIUM and LARGE settings. However, machine learning-based cost estimation significantly outperforms this approach, providing a substantial performance boost, particularly in cost reduction. From this, we conclude that the quality of the cost prediction model used in Cost Guidance is critical to the performance of SCENT.
Table 8: The machine learning-based cost estimation significantly outperforms constant value cost estimation, providing a substantial performance boost, particularly in cost reduction.
model online discovery inference sampling
modes $>$ 8.0 $\uparrow$ scaff. $>$ 8.0 $\uparrow$ reward $\uparrow$ cost $>$ 8.0 $\downarrow$
SMALL w/o Cost Guidance 510 $\pm$26 5413 $\pm$334 7.38 $\pm$0.02 37.1 $\pm$1.0
Constant 515 $\pm$9 5259 $\pm$150 7.43 $\pm$0.07 45.7 $\pm$2.8
ML predicted 478 $\pm$11 5150 $\pm$87 7.43 $\pm$0.02 23.7 $\pm$4.0
MEDIUM w/o Cost Guidance 9310 $\pm$863 11478 $\pm$823 7.31 $\pm$0.09 1463 $\pm$62
Constant 13121 $\pm$1073 25487 $\pm$907 7.41 $\pm$0.02 1368 $\pm$149
ML predicted 17705 $\pm$4224 52340 $\pm$4303 7.74 $\pm$0.04 1163 $\pm$147
LARGE w/o Cost Guidance 7171 $\pm$291 8767 $\pm$429 7.13 $\pm$0.12 1678 $\pm$63
Constant 12218 $\pm$1725 26348 $\pm$5397 7.44 $\pm$0.12 1369 $\pm$127
ML predicted 11730 $\pm$1777 88592 $\pm$16256 7.84 $\pm$0.06 1014 $\pm$137
G.2.2 Comparison of Cost-Reducing Methods
In this section, we compare our cost-guided SCENT with other methods to reduce the synthesis cost. The first method is to use the forward synthesis cost as a reward function component (we denote this approach as the "cost reward"). The second method that we tested is trained on subsets of the building block library $M$ with the top $n\%$ of the most expensive molecules being removed. While both approaches can effectively reduce the cost of proposed molecules, they lag behind cost-guided SCENT in terms of online discovery metrics, as can be seen in Figure 10. This suggests that cost biasing strikes the perfect balance between cost-effectiveness and diversity as it is able to consistently find more modes.
Figure 10: Plot comparing different approaches to reduce the cost of the molecules in MEDIUM setting on sEH proxy. We observe that Cost Guidance discovers by far the largest number of (high-reward) modes and scaffolds while reducing the cost.
G.3 Dynamic Library Ablation
Figure 11 illustrates how the performance of the reward-based Dynamic Library scales with the size of the initial set of building blocks. The improvement over the baseline (SCENT (C+P)) is substantial when the initial library is small but diminishes as its size increases. We also conduct an ablation study on the library-building mechanism, comparing our method—selecting the top $L$ molecules from the reward-sorted buffer—to a variant where $L$ molecules are randomly sampled from the buffer. Our results indicate that the selection mechanism plays a crucial role and is the primary driver of performance improvements.
Figure 11: Plot comparing the relative improvement in performance between using a reward-based Dynamic Library (SCENT (C+P+D)) as opposed to a uniform-based one over the baseline (SCENT (C+P)).
G.4 Exploitation Penalty vs Other Exploration Techniques
In this section, we compare our Exploitation Penalty to the widely used $\epsilon$-greedy exploration where $\epsilon=0.05$. Additionally, we "decay" variants that linearly decrease the exploration parameters ($\epsilon$ for $\epsilon$-greedy, and $\gamma$ for exploitation penalty) so that they reach 00 at the $1000$-th iteration. Figure 12 compares the exploration techniques on three variations of SCENT: with/without Cost Guidance, and with Dynamic Library. We observe that exploitation penalty with linear decaying (penalty decaying) helps our cost guided SCENT the most, successfully mitigating its exploitative nature. Dynamic Library also benefits from using an exploitation penalty, possibly because the penalty explicitly encourages the model to visit newly added molecules. Table 9 shows that our Exploration Penalty helps the cost-guided SCENT in exploration, especially when accompanied with Dynamic Library.
Figure 12: Comparison of the exploration techniques on three variations of SCENT: with/without Cost Guidance, and with Dynamic Library. We observe that exploitation penalty with linear decaying (penalty decaying) helps our cost guided SCENT while Dynamic Library benefits significantly from using a non-decaying penalty.
Table 9: Comparison of different exploration techniques for sEH in SMALL setting.
SMALL setting
model online discovery inference sampling
modes $>$ 8.0 $\uparrow$ scaff. $>$ 8.0 $\uparrow$ reward $\uparrow$ cost $>$ 8.0 $\downarrow$
w/o cost guidance disabled 613 $\pm$20 7599 $\pm$233 7.42 $\pm$0.05 39.3 $\pm$3.9
uniform 510 $\pm$26 5413 $\pm$334 7.38 $\pm$0.02 37.1 $\pm$1.0
uniform decaying 589 $\pm$12 6880 $\pm$386 7.41 $\pm$0.06 33.2 $\pm$3.9
penalty 521 $\pm$15 3660 $\pm$99 7.47 $\pm$0.02 39.6 $\pm$2.1
penalty decaying 615 $\pm$16 7603 $\pm$339 7.38 $\pm$0.02 39.4 $\pm$2.4
cost guidance disabled 509 $\pm$24 5794 $\pm$337 7.4 $\pm$0.08 23.9 $\pm$7.6
uniform 478 $\pm$11 5150 $\pm$87 7.43 $\pm$0.02 23.7 $\pm$4.0
uniform decaying 560 $\pm$3 6487 $\pm$299 7.43 $\pm$0.04 21.6 $\pm$1.4
penalty 531 $\pm$31 4162 $\pm$279 7.62 $\pm$0.01 26.3 $\pm$4.0
penalty decaying 595 $\pm$20 6839 $\pm$360 7.42 $\pm$0.05 20.7 $\pm$2.6
cost guidance + dyn. library disabled 555 $\pm$43 7098 $\pm$770 7.56 $\pm$0.04 20.5 $\pm$2.6
uniform 596 $\pm$19 7148 $\pm$497 7.56 $\pm$0.05 19.7 $\pm$2.0
uniform decaying 599 $\pm$19 7670 $\pm$334 7.49 $\pm$0.07 21.0 $\pm$0.9
penalty 715 $\pm$15 8768 $\pm$363 7.66 $\pm$0.03 20.7 $\pm$3.0
penalty decaying 639 $\pm$8 8323 $\pm$276 7.66 $\pm$0.01 21.0 $\pm$1.2
Table 10: Comparison of different approaches of reducing the cost of sampled molecules. We observe that all the methods can reduce the average synthesis cost with respect to SCENT without Cost Guidance (w/o C). However, our cost guided SCENT (C) greatly improves the number of discovered modes and scaffolds in MEDIUM and LARGE settings.
model online discovery inference sampling
modes $>$ 8.0 $\uparrow$ scaff. $>$ 8.0 $\uparrow$ reward $\uparrow$ cost $>$ 8.0 $\downarrow$
SMALL SCENT (w/o C) 510 $\pm$26 5413 $\pm$334 7.38 $\pm$0.02 37.1 $\pm$1.0
SCENT (cost reward) 513 $\pm$13 4926 $\pm$187 7.28 $\pm$0.05 11.9 $\pm$1.0
SCENT (380 cheap.) 526 $\pm$7 5208 $\pm$194 7.36 $\pm$0.04 19.4 $\pm$3.4
SCENT (400 cheap.) 551 $\pm$22 5895 $\pm$233 7.4 $\pm$0.06 35.4 $\pm$1.9
SCENT (C) 478 $\pm$11 5150 $\pm$87 7.43 $\pm$0.02 23.7 $\pm$4.0
MEDIUM SCENT (w/o C) 9310 $\pm$863 11478 $\pm$823 7.31 $\pm$0.09 1463 $\pm$62
SCENT (cost reward) 9624 $\pm$513 11738 $\pm$585 7.18 $\pm$0.05 979 $\pm$46
SCENT (48k cheap.) 11509 $\pm$801 14547 $\pm$924 7.25 $\pm$0.1 1275 $\pm$68
SCENT (56k cheap.) 10253 $\pm$793 12939 $\pm$1020 7.21 $\pm$0.13 1359 $\pm$48
SCENT (C) 17705 $\pm$4224 52340 $\pm$4303 7.74 $\pm$0.04 1163 $\pm$147
LARGE SCENT (w/o C) 7171 $\pm$291 8767 $\pm$429 7.13 $\pm$0.12 1678 $\pm$63
SCENT (cost reward) 7447 $\pm$370 9233 $\pm$700 7.2 $\pm$0.04 1075 $\pm$31
SCENT (96k cheap.) 7876 $\pm$592 9632 $\pm$791 7.24 $\pm$0.02 1343 $\pm$6
SCENT (112k cheap.) 6879 $\pm$358 8189 $\pm$342 7.16 $\pm$0.08 1459 $\pm$102
SCENT (C) 12375 $\pm$264 36930 $\pm$5455 7.52 $\pm$0.09 1267 $\pm$159